OpenStudy (haleyelizabeth2017):

Each roller under a conveyor belt has a radius of 0.5 meters. The rollers turn at a rate of 30 revolutions per minute. What is the linear velocity of the conveyor belt? Please no direct answers :)

OpenStudy (baru):

how about a formula v=\(r\omega\)

OpenStudy (haleyelizabeth2017):

Where w is the angular velocity..

OpenStudy (haleyelizabeth2017):

So it would be 0.5 times 30rev/min? We'd need to convert...

OpenStudy (baru):

yes exactly, so the trick is to convert 30 rev/ min to radians/second

OpenStudy (haleyelizabeth2017):

Times 2pi

OpenStudy (haleyelizabeth2017):

Like radius times 2pi right?

OpenStudy (haleyelizabeth2017):

wait

OpenStudy (baru):

the formula v=rw requires r to be in meters and w to be in radians/second

OpenStudy (haleyelizabeth2017):

\[\frac{8rev}{sec}\times\frac{2\pi0.5m}{rev}\]

OpenStudy (haleyelizabeth2017):

This is using the conversion method I learned from a friend...

OpenStudy (haleyelizabeth2017):

25.1m/s right?

OpenStudy (baru):

umm...sorry i dont follow that conversion at all....

OpenStudy (haleyelizabeth2017):

Oh...

OpenStudy (baru):

radius is 0.5 meter, its alerady meters, no conversions required

OpenStudy (baru):

but w is in rev/min we need it to be in radian/second

OpenStudy (haleyelizabeth2017):

yeah...

OpenStudy (baru):

so lets focus on w now

OpenStudy (baru):

30 revolution every minute =30 revolutions in 60 seconds

OpenStudy (baru):

can you calculate how many revolutions every second?

OpenStudy (haleyelizabeth2017):

0.5?

OpenStudy (baru):

yes, now we have w in revolutions/ second now we know that \(360^0 =2\pi\) radians and one complete revolution is 360 degrees

OpenStudy (haleyelizabeth2017):

LOL isn't one rev always 2pi?

OpenStudy (baru):

yes that is exactly what i've said^

OpenStudy (haleyelizabeth2017):

I've known that x'D

OpenStudy (baru):

one rev = 2pi =360 degrees

OpenStudy (baru):

so we have w= 0.5 rev every second how many radians every second is that?

OpenStudy (haleyelizabeth2017):

2pi(0.5)/sec...

OpenStudy (baru):

yes exactly, how how much does that become when you simplify

OpenStudy (haleyelizabeth2017):

3.141592654....aka pi xD

OpenStudy (baru):

haha, thats right

OpenStudy (baru):

so all thats left is v=rw

OpenStudy (haleyelizabeth2017):

1.57?

OpenStudy (baru):

yep

OpenStudy (baru):

any doubts?

OpenStudy (haleyelizabeth2017):

Naw, thanks mate

OpenStudy (baru):

sure :)